Need help on Solubility Equilibria and explanation how to do and steps please! :)

1a) What is the molar solubility of Iron(II) carbonate in 0.10M Sodium Carbonate? The solubility product constant (Ksp) is 2.1x10-11

Na2CO3\rightarrow0.20M Na+ and 0.10M CO32-

1b) What is the molar solubility of Silver phosphate (Ksp= 1.8x10-18) in 0.20M silver nitrate?

AgNO3\rightarrow0.20M Ag+ and 0.20M NO3-
Answer #1

Q1a) Let molar solubility of FeCO3 = s

FeCO3 dissociates as : FeCO3 (s) \rightarrow Fe2+ (aq) + CO32- (aq)

initial concentration of Fe2+ = molar solubility of FeCO3 = s

initial concentration of CO32- = (molar solubility of FeCO3) + (concentration from Na2CO3)

initial concentration of CO32- = s + 0.10 M

Ksp FeCO3 = [Fe2+][CO32-]

2.1 x 10-11 = (s) * (s + 0.10 M)

assuming s << 0.10 M

2.1 x 10-11 = (s) * (0.10 M)

s = (2.1 x 10-11) / (0.10 M)

s = 2.1 x 10-10 M

molar solubility of FeCO3 = s = 2.1 x 10-10 M

Q1b) Let molar solubility of Ag3PO4 = s

Ag3PO4 dissociates as : Ag3PO4 (s) \rightarrow 3 Ag+ (aq) + PO43- (aq)

initial concentration Ag+ = 3 * (molar solubility of Ag3PO4) + (concentration from AgNO3)

initial concentration Ag+ = 3 * (s) + 0.20 M

initial concentration PO43- = s

Ksp Ag3PO4 = [Ag+]3[PO43-]

1.8 x 10-18 = (3s + 0.20 M)3 * (s)

assuming 3s << 0.20 M

1.8 x 10-18 = (0.20 M)3 * (s)

s = (1.8 x 10-18) / (0.20 M)3

s = 2.25 x 10-16 M

molar solubility of Ag3PO4 = s = 2.25 x 10-16 M